OpenStudy (anonymous):

2/3-i how would I solve this?

OpenStudy (jdoe0001):

to simplify or rationalize the denominator, you multiply both top and bottom by the conjugate of the denominator so \(\bf \cfrac{2}{3-i}\\ \textit{let us use the conjugate }3+i\textit{ of the denominator}\\ \cfrac{2}{3-i}\times \cfrac{3+i}{3+i}\implies \cfrac{2(3+i)}{(3-i)(3+i)}\\ \quad \\ \textit{keep in mind that } (a-b)(a+b) = (a^2-b^2)\\ \quad \\ \cfrac{2(3+i)}{(3-i)(3+i)}\implies \cfrac{2(3+i)}{3-i^2}\)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (jdoe0001):

wops, shooot... I have one mistake... one sec

OpenStudy (jdoe0001):

\(\bf \cfrac{2(3+i)}{(3-i)(3+i)}\implies \cfrac{2(3+i)}{3^2-i^2}\)

OpenStudy (jdoe0001):

when expanding the bottom, keep in mind that \(\bf i^2 \implies \sqrt{-1} \times \sqrt{-1} \implies \sqrt{(-1)^2}\implies -1\)

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

you have to do foil to the denominator right?

OpenStudy (jdoe0001):

no, notice \(\bf i^2 \implies \sqrt{-1} \times \sqrt{-1} \implies \sqrt{(-1)^2}\implies -1\) that means that \(\bf i^2 = -1\) so \(\bf \cfrac{2(3+i)}{3^2-i^2}\implies \cfrac{6+2i}{9-(-1)} \implies \cfrac{6+2i}{9+1}\)